OpenStudy (anonymous):

Find slope and the y-intercept of each line. y=4 and 2-x=3 Will be given a medal.

OpenStudy (texaschic101):

y = 4 is a horizontal line with a slope of 0, and the y intercept is (0,4) 2 - x = 3 -x = -2 + 3 -x = 1 x = -1 This is a vertical line with an undefined slope. Sorry, but I am not 100% sure, but I do not think it has a y intercept.